OpenStudy (crashonce):

The speed of 3 ships are in the ration A:B:C. If they sail for the same distance, what is the ratio of the taken time.

Parth (parthkohli):

Let the speeds be \(kA, ~kB,~ kC\), and the distance be \(x\).\[t_1 = \dfrac{x}{kA}\]\[t_2 = \dfrac{x}{kB}\]\[t_3 = \dfrac{x}{kC}\]\[\Rightarrow t_1 : t_2: t_3 = ~?\]

OpenStudy (crashonce):

1/a:1/b:1/c

OpenStudy (michele_laino):

If I call with d the same distance, then, I can write: \[d:d:d=At _{a}:Bt _{b}:Ct _{c}\] so: \[At _{a}:Bt _{b}:Ct _{c}=1:1:1\] from the first two ratios, I can write: \[At _{a}:Bt _{b}=1:1\] and now, I apply the fundamental property of proportions, so: \[\frac{ t _{b} }{ t _{a} }=\frac{ 1/B }{ 1/A }\] similarly I get, if I consider the proportion between the second and third ratios

OpenStudy (michele_laino):

namely, i f I consider the proportion between the second and third ratios, I get: \[\frac{ t _{b} }{ t _{c} }=\frac{ 1/B }{ 1/C }\]

OpenStudy (crashonce):

@ParthKohli so is my answer corect

OpenStudy (crashonce):

@ParthKohli

OpenStudy (crashonce):

@Michele_Laino whats the answer cos I cant read ur latex

OpenStudy (crashonce):

@perl a little help please

OpenStudy (michele_laino):

If I call with d the same distance, then, I can write: d:d:d=At_a:Bt_b:Ct_c so: At_a:Bt_b:Ct_c=1:1:1 from the first two ratios, I can write: At_a:Bt_b=1:1 and now, I apply the fundamental property of proportions, so: t_b/t_a=(1/B)/(1/A) similarly I get, if I consider the proportion between the second and third ratios namely, i f I consider the proportion between the second and third ratios, I get: t_b/t_c=(1/B)/(1/C)

OpenStudy (crashonce):

so is the ratio of th time 1/a:1/b:1/c

OpenStudy (crashonce):

@Michele_Laino ?

OpenStudy (michele_laino):

yes!

OpenStudy (crashonce):

finally.. thanks

OpenStudy (michele_laino):

thanks!